# FlexTaF: Enhancing Table Reasoning with Flexible Tabular Formats

### v1

@@ Appendix D Details of Training the Classifier in FlexTaF-Single
In this section, we introduce the details of training the classifier in FlexTaF-Single. Our model is implemented with PyTorch (Paszke et al. 2019) and transformers (Wolf et al. 2020). We present training details in Table 17, and we summarize the training data size employed for training the classifier to predict each LLM across various datasets in Table 18. We count the proportion of each tabular format in the training data for different LLMs and datasets, as shown in Figure 7. Specifically, to compare the proportions of different tabular formats in the training data, we calculate the proportion of each format to all labels of all instances in the training data.
@@ Appendix F Comparison FlexTaF with Oracle
In this section, we compare the performance of FlexTaF with that of Oracle, as shown in Table 19. We observe that: (i) The excellent oracle performance shows that there exist differences between different formats, which further suggests that different instances have their suitable tabular formats. (ii) The performance of FlexTaF-Single is limited by classification and has a gap with the oracle performance, since the Pre-trained Language Model (PLM) cannot predict the behavior of LLMs well (Qin et al. 2024; Liu et al. 2024; Bowman 2023). However, we do not conduct experiments to fine-tune LLMs for classification limited by computing resources. (iii) The performance of FlexTaF-Vote also has a gap with the oracle performance, because the voting mechanism we adopt does not make full use of the rich semantic information in the LLM solutions, causing the limited improvement (Ni et al. 2023).

### v2

@@ Appendix D Details of Training the Classifier in FlexTaF-Single
In this section, we introduce the details of training the classifier in FlexTaF-Single. Our model is implemented with PyTorch Paszke et al. (2019) and transformers Wolf et al. (2020). We present training details in Table 17, and we summarize the training data size employed for training the classifier to predict each LLM across various datasets in Table 18. We count the proportion of each tabular format in the training data for different LLMs and datasets, as shown in Figure 7. Specifically, to compare the proportions of different tabular formats in the training data, we calculate the proportion of each format to all labels of all instances in the training data.
@@ Appendix F Comparison FlexTaF with Oracle
In this section, we compare the performance of FlexTaF with that of Oracle, as shown in Table 19. We observe that: (i) The excellent oracle performance shows that there exist differences between different formats, which further suggests that different instances have their suitable tabular formats. (ii) The performance of FlexTaF-Single is limited by classification and has a gap with the oracle performance, since the Pre-trained Language Model (PLM) cannot predict the behavior of LLMs well Qin et al. (2024); Liu et al. (2024b); Bowman (2023). However, we do not conduct experiments to fine-tune LLMs for classification limited by computing resources. (iii) The performance of FlexTaF-Vote also has a gap with the oracle performance, because the voting mechanism we adopt does not make full use of the rich semantic information in the LLM solutions, causing the limited improvement Ni et al. (2023).